### v1

# Holistic Evaluation of Language Models

Abstract
As language models (LMs) are increasingly adopted across domains, high-quality benchmarking frameworks that accurately estimate performance are essential for guiding deployment decisions. While frameworks such as Holistic Evaluation of Language Models (HELM) enable broad evaluation across tasks, they often rely on fixed prompts that fail to generalize across LMs, yielding unrepresentative performance estimates. Unless we estimate each LM’s ceiling (maximum achievable via changes to the prompt), we risk underestimating performance. Declarative prompting frameworks, such as DSPy, offer a scalable alternative to manual prompt engineering by crafting structured prompts that can be optimized per task. However, such frameworks have not been systematically evaluated across established benchmarks. We present a reproducible DSPy+HELM framework that introduces structured prompting methods which elicit reasoning, enabling more accurate LM benchmarking. Using four prompting methods, we evaluate four frontier LMs across seven benchmarks (general/medical domain) against existing HELM baseline scores. We find that without structured prompting: (i) HELM underestimates LM performance (by 4% average), (ii) performance estimates vary more across benchmarks (++2% standard deviation), (iii) performance gaps are misrepresented (leaderboard rankings flip on $3/7$ benchmarks), and (iv) introducing reasoning (chain-of-thought) reduces LM sensitivity to prompt design (smaller performance $\Delta$ across prompting methods). To our knowledge, this is the first large-scale benchmarking study to empirically characterize LM behavior across benchmarks and prompting methods, showing that scalable performance ceiling estimation enables more decision-useful benchmarks. We open-source (i) DSPy+HELM Integration [^1] DSPy+HELM Integration: https://github.com/stanford-crfm/helm/pull/3893 and (ii) Prompt Optimization Pipeline [^2] Prompt Optimization Pipeline: https://github.com/StanfordMIMI/dspy-helm.
1 Introduction
Language models (LMs) have rapidly advanced in text generation, spurring deployment across diverse domains (thirunavukarasu2023large; van2024adapted; seo2024evaluation). Yet, integrating LMs into downstream workflows remains challenging as LMs frequently commit errors (aali2025medval). Even state-of-the-art general-purpose frontier LMs exhibit non-trivial hallucination rates (wang2024prompt; sivarajkumar2024empirical; bang2025hallulens; tamber2025benchmarking). Such concerns are compounded by LMs’ sensitivity to prompt design (razavi2025benchmarking), introducing variability in leaderboard performance.
While benchmarking frameworks such as Holistic Evaluation of Language Models (HELM) (liang2022holistic; bedi2025medhelm) enable holistic evaluation via a comprehensive suite covering diverse tasks, public leaderboards typically evaluate multiple LMs under a fixed prompt per benchmark. However, fixed prompts rarely generalize well across LMs, leading to unrepresentative performance estimates that obscure underlying strengths and weaknesses of LMs. Hence, broader LM adoption necessitates scalable estimation of performance ceilings (i.e., the maximum achievable via prompt-only changes), thereby allowing practitioners to weigh cost–benefit tradeoffs and choose the right model for each downstream task.
Prompt engineering has emerged as a practical alternative to fine-tuning. Well‑designed prompts improve performance, as demonstrated by nori2024medprompt; maharjan2024openmedlm, combining few-shot selection, chain-of-thought (CoT) (wei2022chainofthought), and ensembling. However, these methods rely on hand-engineered prompts, demanding domain expertise and iterative experimentation, making them labor-intensive and often non-robust to new model rollouts (wang2025perspective). Consequently, researchers have explored automatic prompt optimization (APO) (li2025surveyautomaticpromptengineering), which treats prompt design as an optimization problem.
Figure 1: Pipeline overview. (a) DSPy takes HELM’s baseline prompt and produces structured prompt variants. (b) HELM evaluates models under each prompt variant. With structured prompting, we observe more robust evaluation: (i) improved performance, (ii) reduced variance, (iii) altered gaps (flipped rankings).
DSPy (khattab2023dspy) is a widely used declarative framework that represents prompts as modular, parameterized components with an intuitive structure that allows moving from zero-shot prompts to more adaptive prompting styles all within a single unified system supporting reproducible, structured prompting. Moreover, DSPy supports automatic prompt optimizers (APOs) such as MIPROv2 (opsahl2024optimizing), which can convert high-level task specifications into optimized instructions and few-shot examples.
However, declarative prompting frameworks such as DSPy have not been systematically evaluated across established benchmarks, and there is limited guidance on how structured prompting compares with existing baselines. To address this gap, we integrate DSPy with HELM (Figure 1) and present:
  1. 1.
    A reproducible DSPy+HELM framework that introduces structured prompting methods which elicit reasoning, enabling more robust evaluation of LMs across HELM benchmarks.
  2. 2.
    An evaluation of prompting methods (Zero-Shot, Bootstrap Few-Shot with Random Search, MIPROv2) against HELM’s baseline across four LMs and seven HELM benchmarks that span general and medical domains (reasoning, knowledge QA, problem-solving, error-classification), where each prompting method leverages a distinct mechanism for refining prompts to estimate LM performance ceilings.
  3. 3.
    Empirical evidence that without structured prompting: (i) HELM underestimates LM performance (by 4% average), (ii) performance estimates vary more across benchmarks (++2% standard deviation), (iii) performance gaps are misrepresented (leaderboard rankings flip on $3/7$ benchmarks), and (iv) introducing reasoning (CoT) reduces LM sensitivity to prompt design (smaller $\Delta$ across prompts).
Prompt 1: HELM Baseline Given a patient note and a clinical question, compute the requested medical value. Patient Note and Question: ——————————————————————–
Prompt 2: Zero-Shot CoT Your input fields are: “INPUTS”
Your output fields are: “REASONING” and “OUTPUT”
Your objective is: Given the fields “INPUTS”, produce the fields “OUTPUT”
INPUTS:
Given a patient note and a clinical question, compute the requested medical value.
Patient Note and Question: ——————————————————————–
Respond with the corresponding output fields, starting with “REASONING”, then “OUTPUT”.
Prompt 3: BFRS (Few-Shot Optimized) Your input fields are: “INPUTS”
Your output fields are: “REASONING” and “OUTPUT”
Your objective is: Given the fields “INPUTS”, produce the fields “OUTPUT”
IN-CONTEXT EXAMPLES ($K$ Demos):
INPUTS: <input text> $\rightarrow$ REASONING: <steps>, OUTPUT: <output text>
INPUTS:
Given a patient note and a clinical question, compute the requested medical value.
Patient Note and Question: ——————————————————————–
Respond with the corresponding output fields, starting with “REASONING”, then “OUTPUT”.
Prompt 4: MIPROv2 (Instruction + Few-Shot Optimized) Your input fields are: “INPUTS”
Your output fields are: “REASONING” and “OUTPUT”
Your objective is:
You are a highly skilled medical expert working in a busy emergency room. A patient presents with a complex medical history and concerning symptoms. The attending physician needs your immediate assistance in calculating a critical risk score to guide treatment decisions. The patient’s life may depend on your accuracy.
IN-CONTEXT EXAMPLES ($K$ Demos): INPUTS: <input text> $\rightarrow$ REASONING: <steps>, OUTPUT: <output text>
INPUTS:
Given a patient note and a clinical question, compute the requested medical value.
Patient Note and Question: ——————————————————————–
Respond with the corresponding output fields, starting with “REASONING”, then “OUTPUT”.
Figure 2: Structured prompting methods evaluated in our study (Zero-Shot CoT, BFRS, MIPROv2). Each box corresponds to one method, showing how instructions and context differ across methods. For BFRS and MIPROv2, $K$ denotes the number of in-context demonstrations (Inputs $\rightarrow$ Reasoning, Output).
Model API Identifier Release Context Reasoning
Claude 3.7 Sonnet anthropic/claude-3-7-sonnet-20250219 02/19/2025 200k
Gemini 2.0 Flash google/gemini-2.0-flash-001 02/01/2025 1000k
GPT 4o openai/gpt-4o-2024-05-13 05/13/2024 128k
o3 Mini openai/o3-mini-2025-01-31 01/31/2025 200k
Table 1: Language models evaluated in our study. Columns show API identifiers, release dates, maximum context windows, and native reasoning modes (yes/no). We choose widely used models to evaluate whether prompting meaningfully affects even top-tier frontier models. All experiments were run in August 2025.
2 Methodology
DSPy (khattab2023dspy) is a framework for composing modular LM pipelines. Formally, let $\Phi$ denote a LM program with $m$ modules. Each module $i$ has a prompt template $p_{i}$ containing a set of variables (open slots) for the instruction and $K$ demonstration examples. Let $V$ be the set of all such prompt variables across $\Phi$, and let $V\to S$ denote an assignment of each variable $v\in V$ to a concrete string $s\in S$. We write $\Phi_{V\to S}$ to denote running program $\Phi$ under a particular prompt assignment. Given a dataset D=(x,y)D={(x,y)} of inputs $x$ with ground-truth $y$ and an evaluation metric $\mu$ that compares the program’s output $\Phi(x)$ against $y$, the optimization maximizes $\mu$ over all instructions and demonstrations:
$\Phi^{*}=\operatorname*{arg\,max}_{V\to S}\frac{1}{|D|}\sum_{(x,y)\in D}\mu\Big(\Phi_{V\to S}(x),y\Big).$ (1)
2.1 Prompting Methods
Baseline Prompting
As a baseline, we evaluate LMs using the following prompting methods:
1. HELM Baseline. HELM supports multiple prompting configurations; we adopt the commonly reported fixed, zero-shot (hand-crafted) prompt configuration without CoT as the baseline for comparison.
2. Zero-Shot Predict. DSPy’s Zero-Shot Predict configuration is an unoptimized non-adaptive baseline, which we instantiate with the dspy.Predict module. Each module’s instruction prompt is initialized with the same HELM baseline instruction, without in-context demonstrations (i.e. $K=0$).
Structured Prompting
In addition, we evaluate LMs using the following structured prompting methods (Figure 2):
1. Zero-Shot CoT. DSPy’s Zero-Shot CoT configuration utilizes the same prompting structure as Zero-Shot Predict, but instead instantiates the dspy.ChainOfThought module, which elicits step-by-step rationales, instructing the LM to generate an explicit reasoning trace with the output.
2. BFRS. Bootstrap Few-Shot with Random Search (BFRS) (Algorithm 1) leverages the idea of bootstrapping and random sampling to select the best few-shot demonstrations (fixed instructions) in two phases: (i) Bootstrapping demonstrations: the LM program $\Phi$ is run on a subset of training inputs to gather traces for each module. Whenever the output of $\Phi(x)$ for an example $x$ achieves a sufficiently high score (on metric $\mu$), the input-output pair is taken as a candidate demonstration. (ii) Random few-shot search: Given demonstration pools, BFRS randomly samples sets of $K$ demonstrations per module, inserts them into the module, and evaluates the program on a validation split. After trying $N$ combinations, the program with the highest score is returned (with hyperparameters $K$ and $N$).
3. MIPROv2. MIPROv2 (Algorithm 2) is an optimizer that jointly selects instructions and $K$ few-shot demonstrations via: (i) bootstrapping demos, (ii) grounded instruction proposals from a proposer LM conditioned on dataset summaries, program structure, exemplar demos, and trial history, and (iii) Bayesian search over instruction-demo pairs. It treats each configuration $\mathbf{v}$ as hyperparameters, learns $p(y\mid\mathbf{v})$ from trial outcomes, and steers toward high-scoring regions. For efficiency, candidates are scored on mini-batches of size $B$, with periodic full-dataset $D$ evaluations of top contenders; the best full-data configuration is returned (with hyperparameters instruction text, demo-set, and $K$).
Benchmark Input $\rightarrow$ Output Task Samples
MMLU-Pro Reasoning Question $\rightarrow$ Answer Multi-Task Reasoning 1,000
GPQA Graduate Question $\rightarrow$ Answer Graduate-Level QA 446
GSM8K Math Problem $\rightarrow$ Solution Numeric Problem-Solving 1,000
MedCalc-Bench Patient Note $\rightarrow$ Computed Value Computational Reasoning 1,000
Medec Medical Narrative $\rightarrow$ Errors Error Classification 597
HeadQA Medical Question $\rightarrow$ Answer USMLE-Style QA 1,000
MedBullets Medical Question $\rightarrow$ Answer USMLE-Style QA 308
Table 2: HELM benchmarks (publicly available) evaluated in our study. Columns summarize each benchmark’s input $\to$ output mapping, underlying task type, and number of test samples. The benchmarks span reasoning, knowledge QA, problem-solving, and error-classification tasks across both general and medical domains.
2.2 Benchmarks
We choose seven benchmarks (Table 2) based on (i) public availability, (ii) task diversity (reasoning, knowledge QA, problem-solving, error classification), and (iii) domain coverage (general/medical).
MMLU-Pro. MMLU-Pro (wang2024mmlu) is an enhanced version of MMLU that focuses on more challenging, reasoning-intensive questions. It expands answer choices from four to ten and removes trivial items, providing a more discriminative measure of higher-order reasoning. The metric $\mu$ is exact match.
GPQA. GPQA (rein2024gpqa) is a graduate-level multiple-choice benchmark covering biology, physics, and chemistry to test advanced reasoning. The metric $\mu$ is the fraction of correct answers (exact match).
GSM8K. GSM8K (cobbe2021gsm8k) consists of grade school math word problems designed to evaluate reasoning. The task requires computing a final numeric answer, and the metric $\mu$ is exact match.
MedCalc-Bench. MedCalc-Bench (khandekar2024medcalc) is a medical calculation benchmark, where the input is a patient note and a question asking for a numerical/categorical value. The evaluation metric $\mu$ is exact match for the risk, severity, and diagnosis categories, and a within-range correctness check for others.
Medec. Medec (abacha2024medec) is an error detection and correction benchmark, where each input contains a narrative that may contain factual errors, and the task is to identify/correct these errors. The evaluation metric $\mu$ involves checking how accurately LMs identify whether a note contains an error (binary).
HeadQA. HeadQA (vilares2019head) is a collection of biomedical multiple-choice questions for testing medical knowledge, where questions cover medical knowledge and often resemble medical board exams. The performance metric $\mu$ is exact match between the prediction and the correct option.
MedBullets. MedBullets (medbullets) is a benchmark of USMLE-style medical questions with multiple-choice answers. MedBullets covers broad topics and is designed to reflect the difficulty of medical licensing exams. Like HeadQA, the primary metric $\mu$ is exact match accuracy on the correct answer.
Algorithm 1 BFRS: Bootstrap Few-Shot with Random Search
1:Seed program $\Phi_{\mathrm{seed}}$; train/val sets $D_{\mathrm{tr}},D_{\mathrm{val}}$; threshold $\tau$; demos per module $K_{i}$; trials $R$; minibatch size $B$.
2:Bootstrap: For each $(x,y)\!\in\!D_{\mathrm{tr}}$: run $\Phi_{\mathrm{seed}}$; if $\mu(\Phi_{\mathrm{seed}}(x),y)\!\geq\!\tau$, then for each module $i$ add $\big(u_{i}(x),\Phi^{(i)}_{\mathrm{seed}}(u_{i}(x))\big)$ to $\mathcal{B}_{i}$.
3:Search: For r=1:Rr\!=\!1{:}R:
4:for i=1:mi=1{:}m do
5:  Sample $S_{i}^{(r)}\leftarrow\mathrm{SampleK}(\mathcal{B}_{i},K_{i})$
6:  Let $\mathbf{v}^{(r)}\!\leftarrow\!(I_{1}^{\mathrm{seed}},S^{(r)}_{1},\dots,I_{m}^{\mathrm{seed}},S^{(r)}_{m})$.
7:  Draw minibatch $\mathcal{B}\!\subset\!D_{\mathrm{val}}$ with $|\mathcal{B}|=B$; compute $\widehat{J}_{B}(\mathbf{v}^{(r)})$ by equation 5.
8:Select: $\mathbf{v}^{\star}\!\in\!\arg\max_{r}\widehat{J}_{B}(\mathbf{v}^{(r)})$; optionally re-evaluate $J(\mathbf{v}^{\star})$ on full $D_{\mathrm{val}}$.
9:Return $\mathbf{v}^{\star}$ and the resulting $\Phi_{\mathbf{v}^{\star}}$.
Algorithm 2 MIPROv2: Joint Optimization of Instructions & Demos
1:Train/val sets $D_{\mathrm{tr}},D_{\mathrm{val}}$; candidate sizes $T_{i}$ (instructions), $K_{i}$ (demos per module); minibatch size $B$; escalation period $E$; TPE quantile $\gamma$; trials $T$.
2:Bootstrap demos: Build $\{\mathcal{B}_{i}\}_{i=1}^{m}$ as in equation 3.
3:Propose instructions: For each $i$, sample $\mathcal{I}_{i}=\{I_{i}^{(t)}\}_{t=1}^{T_{i}}$ from proposer LM using task/program-aware context.
4:Initialize history $\mathcal{H}_{0}\!\leftarrow\!\varnothing$; best full-eval $(\mathbf{v}^{\dagger},J^{\dagger})\!\leftarrow\!(\text{seed},0)$.
5:for t=1:Tt=1{:}T do
6:  Fit/update TPE from $\mathcal{H}_{t-1}$ to obtain $\ell,g$ in equation 6.
7:  Acquire candidate:
$\mathbf{v}^{(t)}\in\arg\max_{\mathbf{v}\in\prod_{i}(\mathcal{I}_{i}\times\mathcal{B}_{i}^{K_{i}})}\frac{\ell(\mathbf{v})}{g(\mathbf{v})}.$
8:  Draw minibatch $\mathcal{B}\subset D_{\mathrm{val}}$, $|\mathcal{B}|=B$, score $y^{(t)}=\widehat{J}_{B}(\mathbf{v}^{(t)})$.
9:  Append to history: $\mathcal{H}_{t}\!\leftarrow\!\mathcal{H}_{t-1}\cup\{(\mathbf{v}^{(t)},y^{(t)})\}$.
10:  if $t\bmod E=0$ then
11:   Select top-$K$ by running mean; evaluate each on full $D_{\mathrm{val}}$ to get $J(\cdot)$.
12:   If any $J(\mathbf{v})>J^{\dagger}$ then update $(\mathbf{v}^{\dagger},J^{\dagger})\!\leftarrow\!(\mathbf{v},J(\mathbf{v}))$.
13:Return $\mathbf{v}^{\dagger}$ and $\Phi_{\mathbf{v}^{\dagger}}$.
Prompting Method Claude 3.7 Sonnet Gemini 2.0 Flash GPT 4o o3 Mini
HELM Baseline 64.81% $\pm$ 22.6 61.41% $\pm$ 23.8 61.04% $\pm$ 23.9 70.93% $\pm$ 19.7
[…]
Ceiling - Baseline ($\Delta$) +4.99% +4.80% +4.83% +2.31%
Table 3: HELM leaderboard (macro-averaged over seven benchmarks) across four language models and five prompting methods. Green marks the "ceiling" performance for a model (best value across prompting methods). Entries are reported as the macro-average $\pm$ standard deviation $\sigma$ over seven benchmarks. At each model’s ceiling, structured prompting on average leads to ++4% in accuracy and -2% in $\sigma$ across benchmarks.
2.3 Experimental Setup
Implementation details. We evaluate four frontier LMs (Table 1) to study whether prompting meaningfully impacts the performance of even top-tier LMs. We seed each DSPy program with the baseline instruction for the corresponding benchmark from HELM’s prompt. For BFRS and MIPROv2 optimizers, we follow DSPy’s data separation: the demonstration pool is bootstrapped exclusively from the training split, while candidate prompts are evaluated on a disjoint held-out validation split from the original training partition; neither optimizer ever sees the HELM leaderboard test set. Each benchmark’s loader creates a fixed train/val partition (default 90/10 with the same seed), and we cap both bootstrapped and labeled demonstrations at $K\leq 3$ per module. All final scoring is performed via HELM, so outputs are judged identically regardless of how they were produced. All results reflect single, deterministic runs (temperature = 0), matching HELM’s experimental setup. For HELM baselines, we report HELM’s public leaderboard scores when the setup matches ours: (i) identical LM API version, (ii) zero-shot prompting, and (iii) no CoT reasoning. For benchmarks where the leaderboard setup does not match, we reproduce them with single, deterministic runs.
Metric calculation. To summarize overall gains, we take the mean of the three structured prompting methods (Zero-Shot CoT, BFRS, MIPROv2); for each LM, we first macro-average across benchmarks, and then average the $\Delta$ (absolute % change over baseline) across LMs. The change in variability is reported analogously using the per-model across-benchmark standard deviation, $\sigma$. For each model and prompting method, $\sigma$ is computed over per-benchmark scores (equal weight per benchmark).
Benchmark Prompting Method Claude 3.7 Sonnet Gemini 2.0 Flash GPT 4o o3 Mini
MMLU-Pro HELM Baseline 76.3% $\pm$ 2.7 66.1% $\pm$ 3.0 62.2% $\pm$ 3.0 77.1% $\pm$ 3.1
[…]
Zero-Shot CoT 79.7% $\pm$ 2.5 75.3% $\pm$ 2.7 67.6% $\pm$ 3.0 76.2% $\pm$ 3.1
[…]
GPQA HELM Baseline 57.0% $\pm$ 4.7 53.4% $\pm$ 4.7 45.5% $\pm$ 4.7 57.6% $\pm$ 4.5
Zero-Shot Predict 62.1% $\pm$ 4.5 54.5% $\pm$ 4.7 41.7% $\pm$ 4.5 66.6% $\pm$ 4.3
[…]
GSM8K HELM Baseline 80.5% $\pm$ 2.5 84.0% $\pm$ 2.3 81.1% $\pm$ 2.5 88.6% $\pm$ 2.0
Zero-Shot Predict 83.0% $\pm$ 2.3 77.3% $\pm$ 2.6 84.6% $\pm$ $2 \cellcolorgreen$!2593.6% $\pm$ 1.6
[…]
Table 4: HELM leaderboard (general domain) across four language models and five prompting methods. Green marks the "ceiling" performance for a model (best value across prompting methods). $\uparrow$ and $\downarrow$ indicate a one-step increase or decrease in leaderboard rank, respectively. Entries are reported as mean $\pm$ 95% bootstrap confidence interval. Overall, structured prompting consistently improves the robustness of benchmarks.
Benchmark Prompting Method Claude 3.7 Sonnet Gemini 2.0 Flash GPT 4o o3 Mini
MedCalc-Bench HELM Baseline 21.0% $\pm$ 2.5 15.8% $\pm$ 2.3 18.8% $\pm$ 2.5 34.0% $\pm$ 3.0
Zero-Shot Predict 20.6% $\pm$ 2.6 17.0% $\pm$ 2.4 15.7% $\pm$ 2.3 33.4% $\pm$ 2.9
[…]
MIPROv2 34.7% $\pm$ 3.0 25.4% $\pm$ 2.7 26.8% $\pm$ 2.8 34.3% $\pm$ 3.0
[…]
Zero-Shot Predict 58.3% $\pm$ 3.9 59.3% $\pm$ 4.0 57.3% $\pm$ 4.0 68.3% $\pm$ 3.9
Zero-Shot CoT 61.8% $\pm$ 4.0 59.5% $\pm$ 4.0 59.5% $\pm$ 4.0 68.2% $\pm$ 3.9
[…]
HeadQA HELM Baseline 91.2% $\pm$ 1.8 88.0% $\pm$ 2.1 90.6% $\pm$ 1.8 89.3% $\pm$ 1.9
[…]
MedBullets HELM Baseline 64.9% $\pm$ 5.5 63.0% $\pm$ 5.5 71.1% $\pm$ 5.2 81.2% $\pm$ 4.6
Zero-Shot Predict 65.3% $\pm$ 5.5 64.9% $\pm$ 5.2 71.4% $\pm$ 5.2 81.5% $\pm$ 4.2
Zero-Shot CoT 71.8% $\pm$ $\cellcolorgreen!25$70.8% $\pm$ 5.2 72.1% $\pm$ 5.2 81.5% $\pm$ 4.6
BFRS 71.4% $\pm$ 5.2 69.5% $\pm$ 5.2 72.7% $\pm$ $\cellcolorgreen!25$82.5% $\pm$ 4.6
[…]
Table 5: MedHELM leaderboard (medical domain) across four language models and five prompting methods. Green marks the "ceiling" performance for a model (best value across prompting methods). $\uparrow$ and $\downarrow$ indicate a one-step increase or decrease in leaderboard rank, respectively. Entries are reported as mean $\pm$ 95% bootstrap confidence interval. Overall, structured prompting consistently improves the robustness of benchmarks.
3 Results and Discussion
Figure 3: Heat map showing $\Delta$ (increase in accuracy) of each prompting method over HELM’s baseline (light=small, dark=large). Across four models, x-axis lists prompting methods, y-axis lists benchmarks. All structured prompting methods exhibit similar improvements, while o3 Mini remains relatively insensitive.
3.1 Impact of Structured Prompting on HELM Leaderboard
Improved performance over HELM baseline. Across LMs and benchmarks, structured prompting methods (Zero-Shot CoT, BFRS, MIPROv2) consistently improve over the HELM baseline (Table 3). On average, LMs gain ++4% in absolute accuracy. Non-reasoning models benefit most (++5%), while the reasoning-optimized o3 Mini sees smaller but consistent gains (++2%).
Flipped leaderboard rankings. Taking ceiling performance into account alters three leaderboard rankings. On MMLU-Pro (Table 4), baseline o3 Mini > Claude 3.7 Sonnet (77.1% vs. 76.3%) reverses to Claude 3.7 Sonnet > o3 Mini (80.6% vs. 78.4%). On GSM8K, GPT 4o overtakes Gemini 2.0 Flash, shifting from (81.1% vs. 84.0%) to (90.7% vs. 84.2%). Finally, on MedCalc-Bench (Table 5), baseline o3 Mini > Claude 3.7 Sonnet (34.0% vs. 21.0%) becomes Claude 3.7 Sonnet > o3 Mini (35.3% vs. 34.7%), highlighting how prompting can act as a confounder in leaderboard rankings.
Altered inter-model performance gaps. When evaluated at ceiling performance (best score across prompting methods), models can either narrow or widen their relative performance gaps, providing a more accurate view of true capability differences. Averaging across benchmarks, the gap between the top two models (o3 Mini and Claude 3.7 Sonnet) shrinks from 6% at baseline (70.9 vs. 64.8) to 3% (73.2 vs. 69.8). However, this trend is not uniform: on GPQA, the gap widens substantially, from 0.6% at baseline (57.6 vs. 57.0) to 4.3% at ceiling (68.4 vs. 64.1).
Reduced across-benchmark variance. Structured prompting methods reduce dispersion. Across-benchmark $\sigma$ drops for Claude 3.7 Sonnet (22.6% $\rightarrow$ 18.8%), Gemini 2.0 Flash (23.8% $\rightarrow$ 20.9%), and GPT 4o (23.9% $\rightarrow$ 22.5%), while o3 Mini is unchanged (19.7%), indicating lower sensitivity.
Benchmark-dependent sensitivity. Performance gains vary across benchmarks (Figure 3). Tasks requiring reasoning, such as MMLU-Pro, GPQA, GSM8K, MedCalc-Bench, and MedBullets, show the largest gains (average ++5.5% absolute across models). In contrast, HeadQA and Medec exhibit smaller improvements (average ++0.4% absolute across models). We hypothesize that HeadQA is bottlenecked by high baseline scores ($\sim$90%), while Medec likely reflects fundamental limits in the LM’s knowledge base.
Ranking stability analysis. To assess how structured prompting shifts relative model rankings, we compute mean ranks (1 = best, 4 = worst) across all seven benchmarks. Under structured prompting, the ranking spread compresses: o3 Mini remains the top model but becomes less dominant (1.29 $\to$ 1.57), Claude 3.7 Sonnet improves and moves closer to the top (2.29 $\to$ 2.00), GPT 4o also improves modestly (3.14 $\to$ 3.00), while Gemini 2.0 Flash slightly declines (3.29 $\to$ 3.43). Rank standard deviation ($\sigma$) shows a similar but more moderate pattern: Claude 3.7 Sonnet (0.95 $\to$ 1.15) and GPT 4o (0.90 $\to$) 1.00) become slightly more volatile, while Gemini 2.0 Flash stabilizes (0.76 $\to$ 0.53), and o3 Mini remains roughly unchanged (0.76 $\to$ 0.79). While effects are modest (only 3 instances of ranking flips across leaderboards), the results demonstrate that benchmark rankings are not invariant to prompt design.
3.2 Chain-of-Thought vs Automatic Prompt Optimization
Performance gains primarily attributed to CoT. We study the impact of each prompting method on the leaderboard by averaging results across LMs and benchmarks. Moving from HELM’s baseline to Zero-Shot Predict yields minimal improvement (64.6% $\to$ 64.9%). In contrast, introducing CoT reasoning and moving from Zero-Shot Predict to Zero-Shot CoT results in substantial gains (64.9% $\to$ 68.5%). Interestingly, moving from Zero-Shot CoT to more sophisticated optimizers, such as BFRS and MIPROv2, does not lead to a meaningful additional improvement (68.5% $\to$ 68.6%), indicating that once CoT is introduced, LMs become relatively insensitive to further optimization.
Why CoT reduces sensitivity to prompt design? We provide empirical evidence that reasoning LMs such as o3 Mini, or non-reasoning LMs with prompt-induced reasoning (CoT), are relatively insensitive to prompt design (meincke2025prompting). Now, we provide deeper insight into this phenomenon. With CoT, the model marginalizes over many reasoning paths; prompt optimization mostly reweights sampled paths. A model with parameters $\theta$ reads an input $x$ (a benchmark item) under a prompt $p$, samples a reasoning path $\tau$ (chain-of-thought), and produces a final answer $y$; we compare this to an alternative prompt $p^{\prime}$ (a structured prompt variant). CoT factorizes the joint as: (1) $P_{\theta}(\tau,y\mid x,p)=P_{\theta}(\tau\mid x,p)\,P_{\theta}(y\mid x,\tau,p)$ and predicts by marginalizing over paths (self-consistency): (2) $P_{\theta}(y\mid x,p)=\sum_{\tau}P_{\theta}(\tau\mid x,p)\,P_{\theta}(y\mid x,\tau,p)$. Because (2) is a marginalization of (1), the data-processing inequality implies that the change in the answer distribution between prompts $p$ and $p^{\prime}$ is upper-bounded by the change in the path distribution: $\big\|P_{\theta}(y\mid x,p)-P_{\theta}(y\mid x,p^{\prime})\big\|_{TV}\;\leq\;\big\|P_{\theta}(\tau\mid x,p)-P_{\theta}(\tau\mid x,p^{\prime})\big\|_{TV}\;\leq\;\sqrt{\tfrac{1}{2}\,D_{KL}\!\big(P_{\theta}(\tau\mid x,p)\,\|\,P_{\theta}(\tau\mid x,p^{\prime})\big)}$ where $\|\cdot\|_{\mathrm{TV}}$ is total-variation distance and $\mathrm{KL}(\cdot\|\cdot)$ is Kullback–Leibler divergence (applying the data-processing inequality followed by Pinsker’s inequality). We define the decision margin at item $x$ under prompt $p$ as $m(x;p)=P_{\theta}(y^{\star}\mid x,p)-\max_{y\neq y^{\star}}P_{\theta}(y\mid x,p),\quad y^{\star}=\arg\max_{y}P_{\theta}(y\mid x,p).$ If the left-hand side of the previous bound is $\tfrac{1}{2}\,m(x;p)$, then the prediction is invariant: $\arg\max_{y}P_{\theta}(y\mid x,p^{\prime})\;=\;\arg\max_{y}P_{\theta}(y\mid x,p).$ Since CoT enlarges $m(x;p)$ by averaging diverse valid paths, typical prompt optimization, which mainly reweights $P_{\theta}(\tau\mid x,p)$, cannot flip decisions except on near-tied items, making reasoning models largely prompt-insensitive. If $\mathrm{KL}\!\big(P_{\theta}(\tau\mid x,p)\,\|\,P_{\theta}(\tau\mid x,p^{\prime})\big)\leq\kappa$ and $m(x;p)\geq 2\varepsilon$, then predictions are invariant under $p\!\to\!p^{\prime}$ whenever $\sqrt{\kappa/2}<\varepsilon$.
Figure 4: Accuracy vs cost tradeoff across prompting methods. Each point represents a model-prompt pair, with x-axis showing additional prompt tokens (relative to HELM baseline) and y-axis showing macro-averaged accuracy across benchmarks. Overall, Zero-Shot CoT is the most cost-effective structured prompting method.
3.3 Computational Cost Analysis
We evaluate the inference-time computational cost across prompting methods by examining token usage. BFRS and MIPROv2 optimizations are one-time expenses amortized over future runs: DSPy’s documentation reports that optimization runs range from a few cents to tens of dollars depending on the configuration, [^3] DSPy Optimizer Costs: https://dspy.ai/learn/optimization/optimizers/.. Our optimization costs match DSPy’s reported range. We therefore focus our analysis on inference-time tokens.
In our setup, the input (e.g., question, patient note) is identical across prompting methods, and outputs are capped at <<200 tokens. As a result, differences in inference cost arise almost entirely from the prompt prefix (the instructions and demonstrations prepended to the input). Hence, we quantify the number of additional prompt tokens relative to the HELM baseline instruction, capturing each prompting method’s overhead.
DSPy introduces a lightweight structured prompt template across all methods, resulting in 138 additional tokens for Zero-Shot Predict and 164 tokens for Zero-Shot CoT, which further includes a brief reasoning header. In contrast, BFRS and MIPROv2 insert task-specific demonstrations, producing much larger prompts: averaged across LMs and benchmarks, BFRS adds 1,779 tokens per query and MIPROv2 adds 1,694.
Figure 4 shows the resulting tradeoff. Few-shot optimizers reach high ceilings but require the largest token budgets. Zero-Shot CoT captures most of these gains while using minimal additional prompt tokens, making Zero-Shot CoT the most cost-effective structured prompting method in our study.
4 Related Work
Holistic benchmarking. The General Language Understanding Evaluation (GLUE) (wang2018glue) benchmark was one of the first multi-task evaluation frameworks, aggregating nine distinct language understanding tasks. Benchmarks of increasing scale followed: (i) Measuring Massive Multitask Language Understanding (MMLU) (hendrycks2021measuringmassivemultitasklanguage), including 57 tasks spanning STEM, humanities, social sciences, and (ii) Beyond the Imitation Game (BIG-Bench) (srivastava2023imitationgamequantifyingextrapolating), with 204 diverse tasks. The HELM framework is an established standard, designed for transparent, reproducible, and multi-metric evaluation of model capabilities (liang2022holistic). However, these benchmarks are typically evaluated using static prompts. liang2022holistic note they opt for simple, generic prompts to orient development "towards generic language interfaces" that do not require "model-specific incantations". This reliance on fixed prompts, however, risks the underestimation of the true capabilities of LMs. srivastava2023imitationgamequantifyingextrapolating conclude that standard few-shot prompting "substantially underestimates the best performance and capabilities of language models" when compared to methods that elicit reasoning (suzgun2022challengingbigbenchtaskschainofthought).
Prompting methods. The discovery of in-context learning (brown2020languagemodels), where models learn from n-shot demonstrations, and the breakthrough of chain-of-thought (CoT) prompting (wei2022chainofthought) established the important role of prompt design in model performance. Complex, manually-composed strategies like Medprompt (nori2023medprompt), which combine few-shot selection, CoT, and ensembling, demonstrate that an LM’s performance ceiling often lies higher than with the use of static prompts. Because manual prompt engineering is impractical for systematically finding this ceiling, researchers often frame prompt design as a formal "optimization problem", leading to the field of APO. Early APO methods include generation-and-selection, such as Automatic Prompt Engineer (APE) (zhou2023largelanguagemodelshumanlevel), which uses an LM to propose candidate instructions and a separate scoring function to select the best one. Subsequent systems expanded this search paradigm (wang2023promptagent; yang2023large; singla2024dynamic). These methods often outperform zero-shot or manually engineered prompts on a variety of general tasks. In the LM-as-Optimizer paradigm, an LM is instructed to iteratively refine prompts by showing it a trajectory of previously evaluated candidates and their scores. Other approaches have employed evolutionary search, like Promptbreeder (fernando2024promptbreeder), which treats prompts as "genes" and evolves a population of instructions over generations using a LM to perform mutation. The DSPy framework (khattab2023dspy) generalizes these methods, providing a programming model that compiles declarative, multi-stage pipelines.
5 Limitations
Our study has several limitations. First, we focus on widely used frontier LMs rather than open-source models; while this choice highlights that even strong LMs remain sensitive to prompt design, it limits the generality of our findings. Second, our benchmark suite, though diverse, is not exhaustive, and future work should evaluate structured prompting across additional domains. Third, we evaluate a subset of structured prompting methods from the DSPy-family; alternative frameworks could yield higher ceilings. However, our goal is not to identify the optimal prompting method, but to demonstrate that fixed-prompt (without CoT) leaderboard evaluations systematically underestimate LM performance and often distort model comparisons.
6 Conclusion
By integrating DSPy with HELM, we empirically estimate LM performance ceilings, obtaining more representative estimates. Across seven benchmarks and four LMs, we find that without structured prompting: (i) HELM underestimates LM performance (by 4% average), (ii) performance estimates vary more across benchmarks (++2% standard deviation), (iii) performance gaps are misrepresented (leaderboard rankings flip on $3/7$ benchmarks), and (iv) introducing reasoning (chain-of-thought) reduces LM sensitivity to prompt design (smaller performance $\Delta$ across prompting methods). Sensitivity is heterogeneous: reasoning LMs show marginal gains, whereas some benchmarks for non-reasoning LMs benefit more. Gains are largely agnostic to the structured prompting method. The key driver of improvement is the transition from the baseline prompt to any CoT variant, with Zero-Shot CoT providing the most cost-efficient instantiation. Together, we show that scalable and automated performance ceiling estimation enables more robust, decision-useful benchmarks.
Data Availability
The datasets used in this study are fully open-source, and include: MMLU-Pro (wang2024mmlu), GPQA (rein2024gpqa), GSM8K (cobbe2021gsm8k), MedCalc-Bench (khandekar2024medcalc), Medec (abacha2024medec), HeadQA (vilares2019head), and MedBullets (medbullets). Further distribution is subject to the data sharing agreements stipulated by the original creators.
Acknowledgments
AA is supported by NIH grant R01 HL167974 and ARPA-H contract AY2AX000045. NHS acknowledges support from the Debrah and Mark Leslie endowment for AI in Healthcare, and salary support from Stanford Healthcare. ASC receives research support from NIH grants R01 HL167974, R01HL169345, R01 AR077604, R01 EB002524, R01 AR079431, P41 EB027060; ARPA-H contracts AY2AX000045 and 1AYSAX0000024-01; and NIH contracts 75N92020C00008 and 75N92020C00021. Unrelated to this work, ASC receives research support from GE Healthcare, Philips, Microsoft, Amazon, Google, NVIDIA, Stability; has provided consulting services to Patient Square Capital, Chondrometrics GmbH, and Elucid Bioimaging; is co-founder of Cognita; has equity interest in Cognita, Subtle Medical, LVIS Corp, and Brain Key; EA receives consulting fees from Fourier Health. This research was, in part, funded by the Advanced Research Projects Agency for Health (ARPA-H). The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the United States Government.
Competing Interest
No competing interests to declare.
Appendix A Bootstrap Few-Shot with Random Search (BFRS)
Problem setup.
Let $\Phi$ be a LM program with $m$ modules $\{\Phi_{i}\}_{i=1}^{m}$, each parameterized by (i) an instruction string $I_{i}\in\mathcal{I}_{i}$ and (ii) an ordered list of $K_{i}$ few-shot demonstrations $S_{i}=[(u^{(1)}_{i},v^{(1)}_{i}),\dots,(u^{(K_{i})}_{i},v^{(K_{i})}_{i})$] drawn from a module-specific pool $\mathcal{B}_{i}$. Write the prompt configuration as $\mathbf{v}\!=\!(I_{1},S_{1},\dots,I_{m},S_{m})\in\mathcal{V}$, and define the end-to-end score on $(x,y)$ as $\mu(\Phi_{\mathbf{v}}(x),y)\in[0,1]$. Given train/validation splits $D_{\mathrm{tr}},D_{\mathrm{val}}$, the objective is
$J(\mathbf{v})\;=\;\frac{1}{|D_{\mathrm{val}}|}\!\sum_{(x,y)\in D_{\mathrm{val}}}\!\mu\big(\Phi_{\mathbf{v}}(x),y\big)\,.$ (2)
Bootstrapping demonstration pools.
BFRS constructs candidate demo pools $\{\mathcal{B}_{i}\}_{i=1}^{m}$ via rejection sampling with a seed program $\Phi_{\mathrm{seed}}$ (typically zero-shot):
$\mathcal{B}_{i}\;=\;\Big\{\,(u_{i}(x),\,\hat{v}_{i}(x))\;\Big|\;(x,y)\!\in\!D_{\mathrm{tr}},\;\mu\!\big(\Phi_{\mathrm{seed}}(x),y\big)\!\geq\tau,\;\hat{v}_{i}(x)=\Phi^{(i)}_{\mathrm{seed}}\!\big(u_{i}(x)\big)\Big\},$ (3)
where $u_{i}(x)$ denotes input to module $i$ induced by running $\Phi_{\mathrm{seed}}$ on $x$, and $\tau\!\in\![0,1]$ is an acceptance threshold.
Random search over few-shot sets.
With module instructions fixed at $I_{i}\!=\!I_{i}^{\mathrm{seed}}$, BFRS draws $R$ independent candidates
$S_{i}^{(r)}\sim\textstyle\mathrm{SampleK}\!\big(\mathcal{B}_{i},K_{i}\big)\quad\text{and}\quad\mathbf{v}^{(r)}\!=\!(I_{1}^{\mathrm{seed}},S^{(r)}_{1},\dots,I_{m}^{\mathrm{seed}},S^{(r)}_{m})$, (4)
evaluates $\widehat{J}_{B}(\mathbf{v}^{(r)})$ on a size-$B$ minibatch of $D_{\mathrm{val}}$,
$\widehat{J}_{B}(\mathbf{v})\;=\;\frac{1}{B}\!\sum_{(x,y)\in\mathcal{B}\subset D_{\mathrm{val}}}\!\mu\big(\Phi_{\mathbf{v}}(x),y\big),$ (5)
and returns the best $\mathbf{v}^{\star}\!\in\!\arg\max_{r\in[R]}\widehat{J}_{B}(\mathbf{v}^{(r)})$; optionally, $\mathbf{v}^{\star}$ is re-scored on the full $D_{\mathrm{val}}$. Since $\mu\!\in\![0,1]$, Hoeffding implies $|\widehat{J}_{B}(\mathbf{v})-J(\mathbf{v})|\leq\sqrt{\tfrac{\ln(2/\delta)}{2B}}$ with prob. $\geq 1-\delta$.
Appendix B MIPROv2
Search space and objective.
As in §A, $\mathbf{v}=(I_{1},S_{1},\dots,I_{m},S_{m})$ parameterizes $\Phi$ and $J(\mathbf{v})$ is defined in equation 2. MIPROv2 jointly searches over instructions and few-shot demos by (a) bootstrapping demo candidates $\{\mathcal{B}_{i}\}$, (b) proposing instruction candidates $\{\mathcal{I}_{i}\}$ via a proposer LM, and (c) using Bayesian Optimization (BO) with a Tree-structured Parzen Estimator (TPE) surrogate to choose $\mathbf{v}$.
Initialization (proposal sets).
For each module $i$, construct
[…]
where $I_{i}^{(t)}\sim q_{i}(\cdot\mid\mathrm{ctx}_{i})$ are sampled by the proposer LM given context $\mathrm{ctx}_{i}$ (dataset summary, module signature/code, exemplar demos, and tip prompts).
Bayesian surrogate via TPE.
Maintain a history $\mathcal{H}_{t}=\{(\mathbf{v}^{(s)},y^{(s)})\}_{s=1}^{t}$ of tried configurations and noisy scores $y^{(s)}=\widehat{J}_{B_{s}}(\mathbf{v}^{(s)})$ from minibatches. Let $y^{\star}$ be the $\gamma$-quantile of $\{y^{(s)}\}_{s=1}^{t}$ (e.g., $\gamma\!=\!0.2$). TPE models
$\ell(\mathbf{v})\;=\;p(\mathbf{v}\mid y<y^{\star}),\qquad g(\mathbf{v})\;=\;p(\mathbf{v}\mid y\geq y^{\star}),$ (6)
and proposes $\mathbf{v}$ to maximize the ratio $\ell(\mathbf{v})/g(\mathbf{v})$ (improvement proxy). With categorical choices per module, TPE factorizes $\mathbf{v}$ across modules/slots and estimates equation 6 from $\mathcal{H}_{t}$ by smoothed frequency models.
Noisy validation and escalation.
Each candidate $\mathbf{v}$ is scored on a minibatch $\mathcal{B}$ of size $B$ by equation 5. Every $E$ trials, the current top-$K$ candidates (by posterior mean or running average) are escalated to full-$D_{\mathrm{val}}$ evaluation; the best full-eval configuration $\mathbf{v}^{\dagger}$ to date is tracked and returned at the end. Concentration of $\widehat{J}_{B}$ to $J$ is controlled by $B$ (Hoeffding bound as above).